OpenStudy (anonymous):

what number can be added to both the numerator and denominator of 16/37 to produce a fraction equal to 5/8? can you please go step by step

OpenStudy (anonymous):

\[\frac{16+x}{37+x} = \frac{5}{8}\]cross multiply to get\[8(16+x) = 5(37+x)\]\[128+8x = 185+5x\]\[3x=57\]\[x=19\]

OpenStudy (anonymous):

Setup the problem first: \[(16+x)/(37+x) = 5/8\] Cross multiply: \[8(16+x) = 5(37+x)\] \[128+8x=185+5x\] \[3x=57\] \[x=19\] Proof: \[(16+19)/(37+19)=35/56=5/8\]